### v1

# UniSage: A Unified and Post-Analysis-Aware Sampling for Microservices

Abstract.
Traces and logs are essential for observability and fault diagnosis in modern distributed systems. However, their ever-growing volume introduces substantial storage overhead and complicates troubleshooting. Existing approaches typically adopt a sample-before-analysis paradigm: even when guided by data heuristics, they inevitably discard failure-related information and hinder transparency in diagnosing system behavior. To address this, we introduce UniSage, the first unified framework to sample both traces and logs using a post-analysis-aware paradigm. Instead of discarding data upfront, UniSage first performs lightweight and multi-modal anomaly detection and root cause analysis (RCA) on the complete data stream. This process yields fine-grained, service-level diagnostic insights that guide a dual-pillar sampling strategy for handling both normal and anomalous scenarios: an analysis-guided sampler prioritizes data implicated by RCA, while an edge-case-based sampler ensures rare but critical behaviors are captured. Together, these pillars ensure comprehensive coverage of critical signals without excessive redundancy. Extensive experiments demonstrate that UniSage significantly outperforms state-of-the-art baselines. At a 2.5% sampling rate, it captures 56.5% of critical traces and 96.25% of relevant logs, while improving the accuracy (AC@1) of downstream root cause analysis by 42.45%. Furthermore, its efficient pipeline processes 10 minutes of telemetry data in under 5 seconds, demonstrating its practicality for production environments.
* Pinjia He is the corresponding author.
1. Introduction
Over the past decade, microservice architectures, which decompose monolithic applications into loosely coupled, independently deployable services, have become foundational to modern software systems (Ali21Survey, ; huang2021sieve, ; steam, ). This architectural paradigm offers significant benefits, such as enhanced scalability and development agility, and has been widely adopted by industry leaders, including Amazon, Google, and Netflix (practice2017microservices, ).
However, as these systems grow in scale and complexity, understanding their behavior becomes increasingly challenging. Observability data (e.g., traces, logs, and metrics) play a vital role in addressing this challenge. Distributed traces record the end-to-end execution of individual requests across components, revealing causal relationships and latency bottlenecks (dapper, ; liu2023prism, ). Logs provide detailed, timestamped records of discrete events and are often used to understand control flow and error messages (logsurvey, ; jiang2024large, ). Metrics, as aggregated numerical indicators of system performance (such as CPU usage or request latency), are useful for detecting performance anomalies (Hades, ).
Figure 1. A real example of the end-to-end execution of a user request in a microservice system and its observability data.:
The sheer volume of this observability data, however, presents a significant hurdle, imposing substantial storage and processing overhead that makes retaining all data impractical (TraStrainer, ). For instance, a large-scale e-commerce system at Alibaba generates approximately 20 petabytes of traces daily (mint, ). While data compression can alleviate storage costs, it fails to address the inherent imbalance in observability data: the vast majority of recorded system behaviors are normal and repetitive, offering limited diagnostic value. Conversely, edge cases and failure indicators, though rare, are critically important. This disparity motivates sampling techniques, which selectively retain a representative and informative data subset, prioritizing anomalous or diverse behaviors, to reduce overhead while preserving analytical value (sifter, ; tracemesh, ; TraStrainer, ; steam, ; mint, ).
Among these observability data types, trace sampling has been extensively studied because traces record fine-grained request execution paths, making them both highly valuable for debugging and extremely costly to store in full. In contrast, metrics are inherently low-dimensional and aggregated over time intervals, and therefore typically do not require sampling (micrObser, ). Logs are rich in semantics and widely used in failure diagnosis; they also incur substantial storage overhead (logshrink, ; logreducer, ). In practice, many logs are repetitive or carry limited diagnostic value, suggesting the need for selective retention. Despite this, systematic approaches to log sampling remain, to the best of our knowledge, have not been well explored.
Current trace sampling strategies are primarily categorized as head-based or tail-based approaches. As illustrated in Figure 1, a trace begins when a user request enters the system and spans are generated as the request propagates through various services. Once all spans are collected, the trace is finalized and is either discarded or stored, depending on the sampling decision. Head-based sampling makes this decision at the start of this lifecycle. State-of-the-art frameworks like Jaeger (Jaeger, ) and Zipkin (Zipkin, ) employ this method, typically using low, random sampling rates (e.g., 0.1%) due to the unpredictability of failures, which often results in missing critical problematic traces. Conversely, tail-based sampling decides at the trace’s completion, allowing for more informed choices based on the full trace information. While numerous tail-based methods aim to reduce storage (huang2021sieve, ; sifter, ; TraStrainer, ; tracemesh, ; TracePicker, ), they predominantly follow a sample-before-analysis paradigm: data is sampled first, and only the retained subset undergoes downstream analysis (Figure 2 (left)). This sequence risks discarding crucial failure symptoms before any diagnosis, thereby limiting the effectiveness of troubleshooting.
Figure 2. Two sampling paradigms: Prior work vs. UniSage.
To overcome these limitations, we advocate for a post-analysis sampling strategy (Figure 2 (right)) that reverses the conventional sample-before-analysis paradigm. Instead of discarding data blindly or heuristically at the outset, this strategy first applies lightweight anomaly detection on the complete data stream and, when anomalies are indicated, performs root cause analysis (RCA). Only after this preliminary diagnostic step does the system selectively preserve a minimal yet highly informative subset of traces and logs for long-term storage or detailed inspection. As shown in Figure 2, this design brings several advantages: it takes anomaly relevance into consideration, leverages the full observability stream to improve diagnostic accuracy, and grounds sampling decisions in explicit analysis results, thereby offering interpretability and transparency.
However, despite its promise, we identify three requirements in implementing post-analysis-aware sampling in microservice systems: (1) it must remain lightweight for production use, (2) achieve sufficiently accurate guidance by leveraging diverse signals, and (3) flexibly support both anomalous and normal scenarios. To this end, we introduce UniSage, a unified and post-analysis-aware sampling framework for microservice systems. For the efficiency and scalability requirement, we design a lightweight analysis pipeline: RCA is triggered only when anomalies are detected, and the RCA itself is based on statistical methods rather than model training. This makes the framework easy to use, model-agnostic, and decoupled from the sampling component; any RCA approach that outputs service-level anomaly scores can be seamlessly substituted. For the diagnostic accuracy requirement, UniSage integrates signals from three modalities (i.e., traces, logs, and metrics) to improve analysis robustness and broaden fault coverage. Moreover, even in the presence of false positives, the impact on sampling remains limited, since only a small fraction of additional data would be retained. Finally, to handle both anomalous and normal scenarios in a unified manner, we propose a dual-pillar sampling strategy: the edge-case-based sampler acts as a safeguard by focusing on topological and behavioral deviations, while the analysis-guided sampler leverages RCA results to retain anomaly-relevant data. Together, these pillars ensure comprehensive coverage of critical signals without excessive redundancy.
To evaluate the effectiveness and efficiency of UniSage, we conduct experiments on two public datasets (Nezha, ) derived from the widely-used TrainTicket (traintickt, ) and Online Boutique (OnlineBoutique, ) benchmarks. Results show that UniSage outperforms state-of-the-art baselines across efficiency, scalability, and diagnostic accuracy. In terms of sampling quality, it achieves 56.5% trace and 96.25% log coverage at only a 2.5% sampling rate, ensuring that anomaly-related and rare data are effectively preserved. For RCA accuracy, UniSage boosts $AC@1$ by 42.45% over the best baseline on average. On efficiency and scalability, the entire pipeline processes 10 minutes of telemetry data within 5 seconds, and reduces total runtime on large-scale datasets by 36.3% compared to the best baseline, demonstrating strong practicality for production deployment.
The key contributions of this paper are as follows:
  • We propose a post-analysis-aware sampling paradigm that overcomes the limitations of traditional sample-before-analysis approaches. Instead of discarding data blindly or heuristically, it incorporates analysis results into sampling decisions to retain the most valuable data for long-term storage.
  • We design and implement UniSage, the first unified analysis and sampling framework for both traces and logs. It features an efficient anomaly analysis module and a dual-pillar sampling strategy, effectively balancing diagnostic accuracy, efficiency, and sampling coverage across anomalous and normal scenarios.
  • We conduct extensive experiments on two widely-used microservice benchmarks, demonstrating that UniSage identifies more valuable traces and logs under the same budget and improves downstream root cause analysis performance compared to state-of-the-art methods.
2. Background and Motivation
2.1. Background
2.1.1. Observability
Originally introduced in control theory (kalman1960general, ), observability refers to the extent to which a system’s internal states can be inferred from its external outputs. In modern software systems, this concept has been adapted to describe the ability to understand system behavior through telemetry data, including metrics, logs, and traces (opentelemetry, ). These signals are commonly referred to as the “three pillars” of observability (ObservabilityBook, ), collectively enabling engineers to monitor complex distributed systems (yu2024monitorassistant, ), diagnose failures (huang2024demystifying, ; jiang2025l4, ), and ensure reliable operation.
Traces records the execution process of a user request, which typically involves multiple operations across different services. Each operation is represented as a span, and a complete trace is composed of one or more spans organized as a directed acyclic graph (DAG), where edges capture the parent-child (caller-callee) relationships between spans. As shown in Figure 1, a user request flows through multiple microservices, each of which generates observability data such as traces, logs, and metrics during execution. These signals are collected and stored in a collector for subsequent analysis. The trace representation of this request is illustrated in the right part of Figure 1. Each span records metadata such as the service and operation (i.e., span name), and duration.
Metrics represent time series data points that track key performance indicators (KPIs), such as system error rate, CPU utilization, and request rate for a given service. The primary objective of metrics is to provide insights into the health and operational status of a specific component or system. Owing to their compact nature, metrics can be efficiently collected at scale, which is particularly advantageous in large, distributed environments. Furthermore, metrics are well-suited for pre-aggregation at the component level prior to collection, thereby minimizing the computational overhead associated with processing and storing (micrObser, ).
Logs are runtime messages generated by services or components to record system events, often categorized by verbosity levels such as info, warn, and error. Unlike traces, logs are not inherently linked to specific user requests or transactions, but have long served as a primary tool for developers and operators to understand system behavior. In modern observability setups, however, logs generated during the lifecycle of a request are often instrumented with the corresponding trace_idtrace\_id, enabling direct correlation between logs and traces (e.g., the error log shown at the bottom of Figure 1 is associated with the trace illustrated above). Due to the large volume of logs generated in modern systems, log compression techniques have been increasingly adopted to mitigate storage overhead (yu2024logCompression, ; logshrink, ; liu2019logzip, ).
2.1.2. Root Cause Analysis
Root Cause Analysis (RCA), also referred to as troubleshooting, is the process of identifying the underlying causes of failures in a software system, such as service outages or performance degradations. It is considered the primary task of observability data, and many efforts have been devoted to identifying root causes in microservice systems (rcasurvey, ; Eadro, ; Nezha, ; PDiagnose, ; microrca, ; microcause, ; microhecl, ). During incident response, on-call engineers must collect and analyze telemetry data to understand the sequence of events leading to a failure. Furthermore, failures can propagate across service dependencies (e.g., a payment service depending on a database) or deployment configurations (e.g., co-located containers on a shared host), often resulting in cascading effects that obscure the true origin of the problem.
2.2. Motivation
Figure 3. The comparison of siloed and unified sampling.
Motivation 1: Lack of a unified sampling strategy across observability data. Modern observability relies on three key telemetry signals: metrics, traces, and logs (kalman1960general, ; opentelemetry, ). A fundamental limitation in current pipelines is the siloed approach to sampling these signals. While trace sampling has been extensively studied (huang2021sieve, ; sifter, ; tracemesh, ; steam, ) and metrics are naturally optimized through aggregation (micrObser, ), logs are typically handled by independent mechanisms that focus on compression or coarse filtering rather than diagnostic context (yu2024logCompression, ; logshrink, ; liu2019logzip, ). This decoupling of sampling decisions creates a critical gap for troubleshooting: the loss of correlated context across modalities.
As shown in Figure 3 (a), when trace and log sampling are decoupled, the retained traces and logs may lack contextual linkage. For example, an anomalous trace capturing a service’s latency spike may be preserved, but the corresponding WARN or INFO logs, which indicate issues such as a depleted connection pool or frequent garbage collection, are discarded by the log system under heavy load. The engineer is left with a clear symptom (the anomalous trace) but deprived of explanatory evidence (the related logs), breaking the diagnostic narrative and delaying resolution. This motivates us to explore a unified sampling framework for traces and logs. By considering their cross-modal correlations, as illustrated in Figure 3 (b), we can retain a coherent and informative subset of data that better supports troubleshooting and mitigation.
Motivation 2: Balancing trade-off between troubleshooting accuracy and storage overhead. Observability data is essential for effective troubleshooting in distributed systems, yet its continuous collection incurs significant storage and processing overhead. Under the prevailing paradigm, where sampling is performed prior to any analysis, there exists an inherent trade-off between data reduction and diagnostic utility. If the sampling process fails to retain anomaly-relevant data points, the accuracy of downstream analysis and troubleshooting inevitably suffers. To address this, our motivation is to invert the conventional order by adopting a post-analysis-aware paradigm. It first performs lightweight anomaly detection on the complete data stream and, when anomalies are indicated, performs RCA. Sampling decisions are then guided by the diagnostic results, allowing microservice systems to selectively retain a minimal yet informative subset of traces and logs that are both anomaly-relevant and edge-representative, thereby mitigating storage overhead without compromising utility.
Motivation 3: Improving interpretability of sampling decisions. Most existing sampling methods operate as black boxes, providing little transparency into why particular traces or logs are retained or discarded. Typically driven by heuristics or learned models, these approaches lack human-understandable justifications, which undermines trust, complicates troubleshooting, and restricts operators’ ability to tune or audit the process. In failure diagnosis, however, understanding the rationale behind preserved (or discarded) data can be as critical as the data itself. To address this, we advocate incorporating interpretable signals that explicitly explain each sampling decision. These signals, such as multi-modal patterns illustrated in Table 1, capture fine-grained behavioral deviations and can serve as direct evidence for diagnosis.
3. Problem Formulation
Consider a large-scale distributed system with $N$ services, where telemetry data (i.e., logs $\mathcal{L}$, metrics $\mathcal{M}$, and traces $\mathcal{T}$) are collected individually for each service. Let $\mathbf{X}$ denote the complete set of telemetry data, with $\mathbf{X}_{\mathcal{F}}$ collected during the fault-free construction phase $\mathcal{F}$ and $\mathbf{X}_{\mathcal{P}}$ collected during the fault-suffering production phase $\mathcal{P}$. The objective is twofold: (1) perform lightweight and accurate fault diagnosis over $\mathbf{X}_{\mathcal{P}}$ using full observability data, and (2) derive a sampling strategy that retains only a small yet informative subset of data for long-term storage and further analysis.
4. Methodology
4.1. Overview
Figure 4 presents the overview of UniSage, a unified post-analysis sampling framework for distributed microservice systems. Unlike conventional approaches that sample data prior to any analysis, UniSage inverts this process by first performing lightweight anomaly detection over multi-modal telemetry streams. By default, UniSage samples edge-representative data to preserve coverage, and only when anomalies are detected does it additionally retain anomaly-relevant data, ensuring sample results are both anomaly-relevant and edge-representative. Specifically, UniSage begins with (1) Unified Data Fusion and Correlation (Section 4.2), where raw telemetry streams are pre-processed and correlated. This is followed by (2) Multi-Modal Anomaly Analysis (Section 4.3), which operates a lightweight model-wise anomaly detection and scores suspicious services if anomalous are detected. Finally, the core of our framework, (3) Dual-Pillar Sampling Strategy (Section 4.4), makes the final sample decision. This strategy is built upon two complementary pillars: an Edge-case-based Sampler that independently seeks out rare and unusual data points, and an Analysis-guided Sampler that prioritizes data confirmed to be relevant to detected issues.
Figure 4. Overview of UniSage.
4.2. Unified Data Fusion and Correlation
The foundational phase of the UniSage is to transform raw, disconnected telemetry streams into a unified and contextually rich representation. This phase is crucial as it establishes the instance-level correlations between traces and logs, enabling the sophisticated cross-modal analysis that follows. The process involves two main stages: initial pre-processing of each telemetry type and the subsequent correlation of traces and logs via their shared trace_idtrace\_id.
4.2.1. Telemetry Pre-processing
To prepare the data for analysis, we apply modality-specific pre-processing techniques to extract structured features from logs, traces, and metrics.
  • Log Parsing: The raw log streams, denoted as a set $\mathcal{L}=\{l_{1},l_{2},\ldots,l_{n}\}$, are typically parsed into a structured format for automated analysis (jiang2024large, ; jiang2024lilac, ). We employ Drain (drain, ), a widely-used online log parsing tool, to convert unstructured log messages into a set of parameterized templates $TP=\{tp_{1},tp_{2},\ldots,tp_{m}\}$. Each log entry $l_{i}$ is thus represented by its template, timestamp, service of origin, and verbosity level (e.g., INFO, WARN, ERROR).
  • Trace Encoding: Distributed traces, which record the end-to-end execution of requests, are inherently structured as directed acyclic graphs (DAGs). Each trace $t\in\mathcal{T}$ is composed of spans, where each span represents an operation within a service. To facilitate topological analysis, we decompose each trace graph $G_{t}$ into a set of unique root-to-leaf call paths, $\mathcal{C}_{t}=\{c_{1},c_{2},\ldots,c_{k}\}$. An individual call path $c_{j}$ is an ordered sequence of operations $O$, where each call path is an ordered sequence of operations $(s,o)$.
  • Metric Aggregating: Metrics are collected as multivariate time series data from each service. These metrics, denoted by $\mathcal{M}=\{m_{1},m_{2},\ldots,m_{n}\}$, are aggregated into fixed-time intervals (e.g., one minute) to form vectors that represent the service’s state at discrete points in time.
4.2.2. Cross-Modal Context Correlation
We assume a modern observability setup where logs generated during the lifecycle of a request are instrumented with the corresponding trace_idtrace\_id. This shared identifier is the key to our unified analysis. Before the main analysis pipeline, we perform a stream-processing step that establishes a direct link between each trace and its associated logs. Formally, for any given trace_idtrace\_id, denoted $t^{id}$, we define a correlation function $C$ that maps the identifier to a contextual tuple: $C(t^{id}\to(t,\mathcal{L}_{t})$, where $t$ is the unique trace corresponding to $t^{id}$, and $\mathcal{L}_{t}$ is the set of all log messages that share the same identifier. This creates a transient mapping for each active request, linking its complete distributed trace with the full set of logs generated during its execution across all services. This unified and correlated context is fundamental for the subsequent cross-modal signal propagation bias (Section 4.4.2).
4.3. Multi-modal Anomaly Analysis
A core insight of the UniSage framework is its post-analysis-aware paradigm, where sampling decisions are guided by a comprehensive analysis performed on the full, unsampled stream of observability data. This section introduces the analysis phase of UniSage, which operates in two stages: a lightweight, low-overhead anomaly detection process, followed by a more in-depth root cause localization that is conditionally triggered only when anomalies are detected. If an anomaly is detected, the outputs of this phase (i.e., service-level anomaly scores) serve as the primary inputs for the subsequent analysis-guided sampler (Section 4.4.2) in our subsequent sampling strategy.
4.3.1. Model-wise Anomaly Detection
The first stage of our analysis is a continuous process of anomaly detection performed independently for each data modality. These detectors are designed to be lightweight, serving as an efficient first pass to identify preliminary signals of system malfunction without incurring significant computational overhead.
  • Log-wise Detection: Building on the parsed log templates $tp_{i}\in TP$ from the pre-processing phase, the log-wise detector focuses on significant shifts in log event frequencies. We primarily monitor the rates of ERROR and WARN level logs per service. We apply the simple yet effective $k$-$\sigma$ rule to detect anomalies. For each service, we compute a weighted log rate combining error and warning logs: $r=w\cdot r^{\text{e}}+(1-w)\cdot r^{\text{w}},$ where $w\in[0,1]$ is a tunable parameter. We set $w$ = 0.8 since error logs are generally more indicative of failures. An anomaly is flagged if this rate deviates from the fault-free mean $\overline{r}$ by more than $k$ standard deviations: $r-\overline{r}>k\cdot\sigma.$
  • Trace-wise Detection: For distributed traces, we use span latency as the primary indicator of performance anomalies. Similar to the log-wise detector, we apply the k-sigma rule to the duration of spans within each service. Services associated with traces containing spans that exhibit abnormal latency are marked as potentially anomalous for further investigation.
[…]
4.3.2. Root Cause Analysis
Once an anomaly is detected, the root cause analysis (RCA) phase further investigates the anomaly and produces a ranked list of root causes, along with interpretable diagnostic reports for SREs to mitigate it. Unlike conventional approaches that sample data before fault analysis, potentially risking the loss of critical information, UniSage conducts RCA on the complete stream of observability data. This design helps balance diagnostic accuracy with storage efficiency. The analysis consists of two steps. First, the Pattern mining identifies deviations from normal behavior and generates pattern events. Each event includes both normal and abnormal values, along with the specific pattern of deviation, allowing SREs to quickly pinpoint anomalous behavior. Then, the Anomaly service scoring prioritizes suspected root cause services based on the observed anomalies.
Table 1. Multi-modal Pattern Templates and Examples.
Modality Pattern Template Example
Log $(ranking,logTemp,service,normal$ $Freq,observedFreq,deviation)$ $(1,GET/api/v1/consignservice/$ $consigns/\dots 502,frontend,0,4.05,4.05)$
Metric $(ranking,metricName,service,normal$ $Avg,observedAvg,deviation)$ $(1,k8s.pod.cpu.usage,$ $consign,0.015,0.241,381.333)$
Trace $(ranking,spanName,service,normal$ $Lat,observedLat,deviation)$ $(1,Session.mergeconsign.entity$ .consignRecord,consign,0.03,0.56,17.91).consignRecord,consign,0.03,0.56,17.91)
Pattern mining.
To handle the heterogeneity and high volume of multi-modal observability data, we design a lightweight pattern miner to extract informative patterns from logs, traces, and metrics. Instead of processing telemetry data streams directly, the miner transforms them into a unified, structured representation of patterns, as illustrated in Table 1. This design ensures processing efficiency and high recall, enabling real-time monitoring without introducing significant latency. For each modality, the miner quantifies the deviation of current behavior from fault-free reference data:
  • Metric Pattern. We generate metric patterns by quantifying the deviation of their current behavior from a known fault-free reference period. We measure the change between the current distribution $Q$ and the reference distribution $P$ using two indicators: the distributional overlap $\mathcal{O}=\int\min(P(x),Q(x))\,dx$, and the relative mean value shift $\Delta^{\mu}$. These indicators are combined into a composite deviation score $\Delta^{\mathcal{M}}=k\cdot\mathcal{O}+(1-k)\cdot\Delta^{\mu},$ where the weighting factor $k$ is fixed at 0.8. This score provides a unified measure that captures both the stability of the metric distribution and the shift in its central tendency.
  • Trace Pattern. For distributed traces, patterns identify end-to-end requests with significant performance degradation. A deviation score $\Delta^{\mathcal{T}}$ is calculated for each trace, representing the maximum relative latency increase of any of its constituent spans compared to its reference latency: $\Delta^{\mathcal{T}}(t)=\max_{sp_{i}\in t}(\frac{lat(sp_{i})-\widehat{lat}(sp_{i})}{\widehat{lat}(sp_{i})})$.
  • Log Pattern. Log patterns are derived from significant increases in ERROR message frequencies. For each ERROR log template $l$, the deviation score $\Delta^{\mathcal{L}}(l)$ is defined as $\Delta^{\mathcal{L}}(l)=\frac{r(l)-\hat{r}(l)}{\hat{r}(l)+\epsilon},$ where $r(l)$ is the current frequency, $\hat{r}(l)$ is the fault-free reference frequency, and $\epsilon$ is a small constant to ensure stability for unseen templates (where $\hat{r}(l)=0$).
Anomaly service scoring.
Based on the mined patterns, we compute a real-time anomaly score for each service $s$, which reflects its likelihood of being the root cause of an issue. Let $A_{s}^{\mathcal{Z}}$ be the set of patterns of modality $\mathcal{Z}$ associated with service $s$. Each pattern $a\in A_{s}^{\mathcal{Z}}$ has a deviation score $\Delta_{a}^{\mathcal{Z}}$ and a rank $R_{a}^{\mathcal{Z}}$ (used to prioritize more severe deviations). The service-level anomaly score is defined as:
$\varphi_{s}=\sum_{\mathcal{Z}\in\{\mathcal{T},\mathcal{L},\mathcal{M}\}}\frac{|A_{s}^{\mathcal{Z}}|}{|A^{\mathcal{Z}}|}\sum_{a\in A_{s}^{\mathcal{Z}}}\frac{\Delta_{a}^{\mathcal{Z}}}{R_{a}^{\mathcal{Z}}+1},$
where $|A_{s}^{\mathcal{Z}}|$ is the number of patterns of modality $\mathcal{Z}$ for service $s$, and $|A^{\mathcal{Z}}|$ is the total number of patterns of that modality across all services. This produces a ranked list of services, which serve as the input for anomalous service bias to guide the analysis-guided sampler (Section 4.4.2) in the subsequent phase.
4.4. Dual-Pillar Sampling Strategy
Figure 5. Dual-Pillar Sampling Strategy of UniSage.
The final phase of the UniSage framework is to synthesize all available information into a robust and efficient sampling decision. We introduce a dual-pillar strategy that decomposes the sampling logic into two complementary components: a Edge-case-based Sampler that provides a reference of coverage for rare and unusual data points, and an Analysis-guided Sampler that precisely targets data related to detected anomalies. This structure ensures that UniSage is both sensitive to known failure patterns and resilient to novel ones.
4.4.1. Edge-case-based Sampler
This Sampler is perpetually active and serves as the foundational component of our sampling strategy. Its purpose is to identify and assign a higher sampling probability to data points that are inherently rare, based on their intrinsic properties rather than the results of the anomaly analysis. This ensures a continuous level of coverage for novel system behaviors and acts as a safeguard against failures that may not be caught by the initial detectors.
Topological Bias.
This bias captures structural deviations in distributed traces. A reference set of call paths $C_{\mathcal{F}}$ is first constructed from fault-free traces. For an incoming trace $t$, we extract its call paths $C_{\mathcal{P}}$ and assess their dissimilarity to the reference set. Specifically, the similarity between a path $c\in C_{\mathcal{P}}$ and a reference path $c^{\prime}\in C_{\mathcal{F}}$ is measured via the Jaccard similarity: $\text{Sim}(c,c^{\prime})=\frac{|c\cap c^{\prime}|}{|c\cup c^{\prime}|}.$ A score of 1 denotes an identical match, while 0 indicates a complete structural deviation. The topological bias score for trace $t$ is then defined as: $Score_{topo}(t)=1-\max_{c^{\prime}\in C_{\mathcal{F}}}\text{Sim}(c,c^{\prime}).$ To reduce computation, if a path $c$ is identical to a reference path, we directly assign $\text{Sim}(c,c^{\prime})=1$ and omit further comparisons. This pruning strategy substantially improves efficiency while preserving accuracy. For interpretability, when $Score_{topo}(t)>0$, the path $c$ and its closest reference $c^{\prime}$ are reported as evidence of structural novelty.
Behavioral Bias.
This bias evaluates the statistical rarity of observed behaviors, including span latencies in traces and template frequencies in logs. For Traces, this bias captures performance anomalies reflected in span latencies. For each span $sp_{i}$ within a trace $t$, we compute its standardized deviation (z-score) with respect to the fault-free distribution of span durations: $z(sp_{i})=\frac{\left|lat(sp_{i})-\mu\right|}{\sigma},$ where $\mu$ and $\sigma$ denote the mean and standard deviation under normal conditions. The behavioral bias score of a trace is defined as the maximum span-level deviation: $Score_{b}(t)=\max_{sp_{i}\in t}z(sp_{i}).$ A higher value indicates that the trace contains at least one span with significantly abnormal latency. Finally, the edge-case sampling probability is determined by combining the behavioral and topological bias scores:
(1) $p_{e}(t)=\frac{2}{1+e^{-2\max(tanh(Score_{topo}(t)),Score_{b}(t))}}-1.$
For Logs, this bias quantifies the rarity of log templates. Let $p_{l}^{\prime}$ denote the empirical probability of template $l$ observed during the fault-free phase $\mathcal{F}$. The behavioral bias score of a log is defined as: $Score_{b}(l)=-\log(p_{l}^{\prime}+\epsilon),$ where $\epsilon$ is a small constant for numerical stability (e.g., $10^{-6}$). Templates unseen in $\mathcal{F}$ are treated as maximally rare by setting $p_{l}^{\prime}=0$. The corresponding sampling probability is then normalized using the same scaled sigmoid function: $p_{e}(l)=\frac{2}{1+e^{-2Score_{b}(l)}}-1.$ This approach ensures that both unseen and low-frequency templates are prioritized during sampling.
4.4.2. Analysis-guided Sampler
Analysis-guided sampler is dynamically activated and is guided by the outputs of the Multi-modal anomaly analysis phase. When the system is operating normally, this sampler has minimal influence. However, upon the detection of an anomaly, it becomes the dominant factor in the sampling decision, ensuring that the most diagnostically relevant data is preserved.
Anomalous Service Bias.
This bias incorporates the final output of the root cause analysis (Section 4.3.2). All data instances (traces and logs) originating from or passing through services with a high anomaly score $\varphi_{s}$ are assigned a proportionally higher sampling probability $p_{a}$. This ensures that data surrounding the most likely source of a failure is comprehensively retained.
For Traces. Given a trace $t$, its analysis-guided sampling probability $p_{a}(t)$ is determined by the maximum anomaly score among all services on its execution path. This is because the most significant anomaly often originates from a specific root cause service. We use the final service score $\varphi_{s}$ from the RCA stage. The probability is defined as: $p_{a}(t)=\frac{2}{1+e^{-2\max_{s\in\mathcal{S}_{t}}\varphi_{s}}}-1,$ where $\mathcal{S}_{t}$ is the set of services involved in trace $t$.
For Logs. Given a log $l$ associated with service $s$, its sampling probability is directly determined by the anomaly score of that service: $p_{a}(l)=\frac{2}{1+e^{-2\varphi_{s}}}-1.$
Cross-Modal Signal Propagation Bias.
This bias leverages the trace–log correlation established in Section 4.2.2 to propagate anomaly evidence across modalities. This ensures that evidence from one signal type can elevate the perceived importance of a correlated signal in another. We update the sampling probabilities $p_{a}(l)$ and $p_{a}(t)$ to reinforce cross-modal diagnostic signals.
  • Trace-to-Log Signal Propagation. The diagnostic value of a log is not solely determined by its content but also by the execution context in which it was generated. A seemingly benign INFO or WARN log can be critically important if it occurred during a highly anomalous transaction. To capture this context, we update the sampling probability of each log $l$ as:
    (2) $p_{a}^{\prime}(l)=\max\big(p_{a}(l),\,1-(1-p_{a}(t_{l}))^{w_{t}}\big),$
    where $t_{l}$ is the trace associated with $l$ and $w_{t}\in(0,1]$ is a weighting factor that controls the propagation strength. This ensures that any log within a highly anomalous trace inherits part of its anomaly likelihood.
  • Log-to-Trace Signal Propagation. Conversely, a trace that appears normal in terms of latency and topology may still indicate a failure if it generates highly anomalous logs. To capture this, we update the sampling probability of a trace $t$ as:
    (3) $p_{a}^{\prime}(t)=\max\big(p_{a}(t),\,1-(1-\max_{l\in\mathcal{L}_{t}}p_{a}(l))^{w_{l}}\big),$
    where $\mathcal{L}_{t}$ is the set of logs associated with $t$ and $w_{l}\in(0,1]$ controls the influence of log signals on the trace. This mechanism ensures that traces leading to explicit errors are prioritized, regardless of their performance characteristics.
4.4.3. Composite Sampler
The final step is to integrate the outputs from the two sampling pillars into a unified decision. The Composite sampler combines the edge-case-based probability $p_{e}$ with the analysis-guided probability $p_{a}$. When the system operates normally and no anomalies are detected, the analysis-guided sampler remains inactive, and the final sampling probability reduces to: $p=p_{e}.$ If anomalies are detected, the analysis-guided sampler is activated, and the final probability is determined by the multiplicative composition: $p=p_{e}\cdot p_{a}.$ This design ensures that the system continues to explore rare or unusual behaviors through $p_{e}$ in normal conditions, while under anomalous conditions, the sampling strategy prioritizes data instances that are both rare and diagnostically relevant. In this way, UniSage balances exploration of novel behaviors with exploitation of anomaly signals.
5. Experiments
This section answers the following research questions:
  • RQ1: How effective is UniSage in trace and log sampling compared to the baseline approaches?
  • RQ2: How effective is UniSage in downstream root cause analysis compared with baseline approaches?
  • RQ3: How efficient and scalable is UniSage in sampling and analysis?
5.1. Datasets
We conduct our experiments on two datasets, TrainTicket (traintickt, ) and OnlineBoutique (OnlineBoutique, ), from Nezha benchmark (Nezha, ). TrainTicket is a railway ticketing microservice system, enabling users to search, book, and pay for train tickets. OnlineBoutique is an e-commerce application, allowing users to browse products, add them to a cart, and complete purchases. Both systems have been widely adopted in prior studies (Nezha, ; Eadro, ; microhecl, ; TraStrainer, ; tracemesh, ), and are deployed on Kubernetes clusters with synthetic workloads simulating user behavior. The datasets contain telemetry data, including traces, metrics, and logs, collected via open-source observability tools (otelcollector, ).
Faults injected into the systems fall into two main categories: resource-related issues (e.g., CPU contention and network jam) and code-level defects (e.g., error returns and exceptions). Specifically, the TrainTicket dataset contains 33 fault injection cases, while the OnlineBoutique dataset includes 56. Additionally, each dataset contains a segment of fault-free data used for reference. We label two types of uncommon traces that are deemed necessary for sampling. The first is abnormal traces collected during the fault injection periods, and the second is edge-case traces that do not appear in fault-free periods. For logs, we label Error-level logs from the services where faults are injected as abnormal. However, some fault cases do not produce any error logs. For such cases, we exclude them from the log-based evaluation. Based on this labeling strategy, the log data contains 9 cases in the OnlineBoutique dataset and 17 in the TrainTicket dataset. Detailed statistics and characteristics of the datasets are summarized in Table 2.
Table 2. Data Statistics.
Datasets # services # traces # spans # logs # metrics
TrainTicket 41 7,049 534,541 201,750 16
OnlineBoutique 11 297,936 13,462,459 3,907,981 16
5.2. Evaluation Metrics
To evaluate the quality of the sampling results, we follow the prior studies (huang2021sieve, ; TraStrainer, ; TracePicker, ; steam, ) to use coverage as an indicator of sampling quality.
  • Coverage measures the extent to which the sampler favors valuable data (i.e., traces or logs). It is defined as the ratio of labeled data selected by the sampler to the total labeled data, where the labels are determined based on fault injection information and manual verification. A higher coverage indicates that the sampler is more effective in retaining informative data for downstream analysis.
To assess the effectiveness of downstream root cause analysis, we employ two commonly used (microrca, ; Eadro, ; Nezha, ) evaluation metrics: Top-k Accuracy ($AC@k$) and Mean reciprocal rank ($MRR$).
  • Top-k Accuracy ($AC@k$) represents the probability that the top $k$ results given by a method include the real root causes. It ranges from 0 to $1$, with higher values indicating better performance. We set $k=1,3$. $AC@k=\frac{1}{|A|}\sum\nolimits_{a\in A}\frac{\sum_{i<k}R^{a}[i]\in V^{a}_{rc}}{min(k,|V^{a}_{rc}|)},$ where $R^{a}[i]$ denotes the $i$th ranking result for the failure case $a$ by an RCA method, and $V^{a}_{rc}$ is the true root cause set of case $a$.
  • Mean reciprocal rank ($MRR$) represents the inverse of the rank of the first identified root cause. If the actual root cause is not present in the result list, the rank is considered to be infinity and its reciprocal is zero. Higher values of MRR indicate better accuracy.
5.3. Baseline Methods
5.3.1. Sampling Baseline
To evaluate the effectiveness and efficiency of UniSage, we compare it against four baseline sampling methods and two Variants of UniSage. Among them, the Uniform strategy is used as a baseline for both trace and log sampling, while the remaining three methods are designed for trace sampling only.
  • Uniform: This strategy uniformly samples a subset of traces without considering any specific characteristics of the traces. It serves as the default sampling mechanism in many distributed tracing systems, commonly referred to as head-based sampling. Since there are no established log sampling methods in prior work, we also adopt this random head sampling approach as the baseline for log sampling.
  • Sieve: Sieve (huang2021sieve, ) is an online tail-based sampling approach, which uses a robust random cut forest (RRCF) (RRCF, ) to detect uncommon traces and sample them with a high probability. Uncommon traces tend to have a shallower depth from the root to leaf, and thus will receive more attention (i.e., a higher probability) for sampling.
  • Sifter: Sifter (sifter, ) captures edge-case traces by learning an unbiased, low-dimensional model based on fixed-length trace sub-paths. Such a model is able to approximate the system’s common-case behaviors. Thus, by measuring the reconstruction loss of an incoming trace, the sampling decisions can be made toward traces that lead to a large loss.
  • TraStrainer: TraStrainer (TraStrainer, ) is a biased sampler that consider not only trace diversity but also system runtime state (i.e., metrics).
  • UniSage w/o $\mathcal{A}$ is a variant that only considers edge cases to set sampling preferences. We achieve this by using the sampling results from the edge-case-based sampler as the final sampling outcome.
  • UniSage w/o $\mathcal{E}$ is a variant that only considers anomalous data to set sampling preferences. We achieve this by using the sampling results from the analysis-guided sampler as the final sampling outcome.
5.3.2. Downstream Analysis Baseline
Root cause analysis (RCA) is a common task performed in downstream analysis. To evaluate the impact of sampling methods on the analysis results, we combined the following three state-of-the-art trace-based RCA methods with different sampling approaches. In addition, to fairly compare our multi-modal RCA approach, we also include PDiagnosis, a widely-used multi-modal RCA method, as a baseline.
  • TraceAnomaly: TraceAnomaly (traceanomaly, ) leverages deep learning to learn normal trace patterns offline, and then detects abnormal traces and identifies root causes online.
  • TraceRCA: TraceRCA (TraceRCA, ) utilizes spectrum analysis to identify root cause services by analyzing the proportion of normal and abnormal invocations.
  • MicroRank: MicroRank (Microrank, ) identifies and ranks root causes by combining personalised Pagerank method and spectrum analysis.
  • PDiagnose: PDiagnose (PDiagnose, ) transforms metrics, traces, and logs into time series and determines root causes through voting abnormal time series.
  • UniSage w/o $\mathcal{M}$ is a variant that drops metric data in anomaly analysis.
  • UniSage w/o $\mathcal{L}$ is a variant that drops log data in anomaly analysis.
5.4. RQ1: Effectiveness of sampling
Figure 6. Trace coverage across sampling budgets for OnlineBoutique (left) and TrainTicket (right).
5.4.1. Trace Sampling
Figure 6 illustrates the coverage performance of various trace sampling methods under different sampling budgets on the OnlineBoutique and TrainTicket datasets, respectively. UniSage consistently achieves the highest coverage across all budget levels on both datasets, demonstrating its effectiveness in capturing failure-related traces. On OnlineBoutique, UniSage attains a coverage of 0.372 at 1% budget, substantially surpassing TraStrainer (0.088), Sieve (0.052), and Sifter (0.003). Even at 10% sampling, UniSage achieves 0.904 coverage, outperforming TraStrainer (0.539) by over 36%. Similarly, on TrainTicket, UniSage reaches 0.839 at just 5% sampling and peaks at 0.975 with 10%, while TraStrainer trails behind with only 0.615 coverage at the same budget. This highlights the robustness of our approach across both datasets.
The Uniform baseline, which performs naive head sampling without considering trace semantics or system context, yields the lowest coverage, often near zero under low sampling budgets. Sifter and Sieve perform slightly better. They aim to approximate normal system behavior and prioritize rare trace types. However, this bias towards underrepresented traces tends to emphasize edge-case behaviors rather than those correlated with actual failures. As a result, their coverage remains relatively low. For example, at 5%, Sieve achieves 0.263 on OnlineBoutique and 0.117 on TrainTicket. TraStrainer, which incorporates runtime signals (i.e., metrics) into its sampling logic, performs notably better than other baselines and consistently ranks second. Nevertheless, its reliance on metric anomalies limits its ability to detect faults that manifest solely in trace or logs. In contrast, UniSage leverages multi-modal signals, including trace, log, and metric data, to detect a broader spectrum of anomaly patterns. This comprehensive perspective enables it to capture anomaly-related traces more accurately and efficiently, resulting in superior coverage across all sampling budgets. Furthermore, the two ablated variants of UniSage both show noticeable performance degradation compared to the full design, confirming that the effectiveness of UniSage critically depends on the joint contributions of its analysis-guided and edge-case-based samplers.
5.4.2. Log Sampling
Table 3. Log Coverage across sampling budgets on two datasets.
Dataset Method 0.1% 1.0% 2.5% 10.0%
OnlineBoutique Uniform 0 0.021 0.037 0.130
UniSage w/o $\mathcal{A}$ 0.347 0.539 0.774 1
UniSage w/o $\mathcal{E}$ 0.574 0.717 0.853 1
UniSage 0.626 0.807 0.943 1
TrainTicket Uniform 0 0.083 0.102 0.129
UniSage w/o $\mathcal{A}$ 0.373 0.945 0.802 1
UniSage w/o $\mathcal{E}$ 0.312 0.797 0.885 1
UniSage 0.373 0.945 0.982 1
Table 3 presents the log coverage achieved by different methods under varying sampling budgets (from 0.1% to 10.0%) on two datasets. The baseline method, Uniform, exhibits limited effectiveness, especially under low budgets. For instance, at 1.0% sampling, Uniform achieves only 2.1% and 8.3% coverage on OnlineBoutique and TrainTicket, respectively. This is expected, as random sampling lacks awareness of log characteristics. In contrast, UniSage demonstrates substantially better performance in all sampling budgets and datasets. Even with just 0.1% of logs, UniSage achieves 62.6% coverage on OnlineBoutique and 37.3% on TrainTicket, significantly outperforming Uniform. As the sampling budget rate increases, UniSage consistently approaches or reaches full coverage, achieving 100% on both datasets at 10.0%. This indicates that UniSage can effectively prioritize problem-related logs. Similarly, the ablation results on log sampling also show a clear performance drop compared to the full UniSage, further confirming the effectiveness of jointly employing both the analysis-guided and edge-case-based samplers.
5.5. RQ2: Effectiveness of root cause analysis
Table 4. Comparison of the impact of different tracing frameworks on the accuracy of root cause analysis.
Paradigm
Tracing
Framework
RCA
Method
Online Boutique Train Ticket
$AC@1$ $AC@3$ MRRMRR $AC@1$ $AC@3$ MRRMRR
Sample-before-Analysis Uniform MicroRank 0.071 0.286 0.206 0.030 0.061 0.065
TraceRCA 0.125 0.268 0.267 0.030 0.091 0.073
TraceAnomaly 0.089 0.268 0.224 0 0.061 0.035
Sifter MicroRank 0.107 0.286 0.244 0.030 0.091 0.089
TraceRCA 0.143 0.304 0.295 0.061 0.121 0.116
TraceAnomaly 0.071 0.250 0.238 0 0.091 0.046
Sieve MicroRank 0.232 0.357 0.381 0.091 0.152 0.175
TraceRCA 0.232 0.321 0.352 0.061 0.152 0.119
TraceAnomaly 0.179 0.268 0.319 0.091 0.182 0.227
TraStrainer MicroRank 0.339 0.411 0.421 0.333 0.394 0.442
TraceRCA 0.286 0.357 0.368 0.273 0.394 0.377
TraceAnomaly 0.304 0.304 0.348 0.242 0.333 0.351
None (Full data) PDiagnosis 0.357 0.446 0.525 0.303 0.424 0.387
UniSage MicroRank 0.357 0.446 0.462 0.394 0.424 0.412
TraceRCA 0.357 0.482 0.497 0.273 0.424 0.399
TraceAnomaly 0.339 0.411 0.386 0.273 0.394 0.357
PDiagnosis 0.339 0.446 0.467 0.303 0.424 0.387
[…]
UniSage 0.411 0.732 0.589 0.545 0.606 0.620
Table 4 reports the downstream performance of various RCA methods (i.e., MicroRank, TraceRCA, TraceAnomaly, and PDiagnosis) under different sampling strategies on two datasets. For fairness, we compare existing approaches strictly within the conventional sample-before-analysis paradigm, where the RCA models operate only on the sampled data. In contrast, our analysis-before-sample approach leverages RCA results derived from the full data stream to guide sampling. This design not only demonstrates the effectiveness of our method but also highlights the superiority of the proposed post-analysis-aware paradigm.
Sampling quality strongly affects RCA effectiveness. From the results under the sample-before-analysis paradigm in Table 4, we observe that higher-quality sampling strategies consistently lead to better RCA accuracy across methods and datasets. The Uniform baseline consistently performs the worst, with $AC@1$ close to zero. Sifter and Sieve show moderate improvements, but their performance remains limited due to their general-purpose design. TraStrainer, which leverages system metrics, performs better and ranks highest among baselines (e.g., $MRR$ of 0.421 on OnlineBoutique with MicroRank), yet it is still lower than UniSage (e.g., $MRR$ of 0.497 on OnlineBoutique with TraceRCA). These results emphasize that sampling strategy quality is critical for effective RCA, especially when the trace budget is constrained. Poorly chosen samples can omit anomaly-relevant data, undermining downstream analysis.
Post-analysis sampling yields superior RCA performance. From the analysis-before-sample paradigm results in Table 4, we observe that UniSage consistently achieves the highest accuracy across all RCA methods and datasets. Unlike baselines that sample first and analyze later, our approach first conducts RCA on the full dataset and then guides sampling with the results, ensuring high coverage of anomaly-relevant data. For instance, with MicroRank, UniSage attains $AC@1$ scores of 0.411 and 0.545 on OnlineBoutique and TrainTicket, respectively, substantially outperforming the best baseline TraStrainer ($AC@1$ of 0.339 and 0.333).
[…]
5.6. RQ3: Efficiency and Scalability
(a) Comparison of sampling efficiency between UniSage and trace-only baselines with Uniform log sampling.
(b) Proportional time consumption of UniSage.
Figure 7. Efficiency and proportional time consumption results.
We assess the runtime efficiency and scalability of different sampling strategies by measuring their total sampling time on both datasets (Table 2). For a fair comparison, we add the time of Uniform log sampling (i.e., random selection) to all trace-only baselines (i.e., Sieve, Sifter, TraStrainer, Uniform), as UniSage uniquely supports joint sampling of both traces and logs. Note that UniSage ’s total time also includes RCA, as it follows a post-analysis sampling strategy, unlike baselines, which measure sampling time only.
As shown in Figure 7(a), UniSage achieves a strong balance between efficiency and scalability. On the large-scale OnlineBoutique dataset (nearly 40x more traces and 19x more logs), UniSage completes end-to-end analysis in 739.4 seconds, significantly faster than Sieve (3,364.1s) and TraStrainer (1,160s), while offering much stronger performance. On the smaller TrainTicket dataset, UniSage finishes in 142.3 seconds (4.3s per case), remaining competitive despite the added RCA overhead. In contrast, some baselines, such as Sifter, show fast runtime on smaller datasets (18.6s), but do not scale well to larger workloads. These results demonstrate that UniSage not only integrates trace and log sampling effectively, but also remains practical for real-world deployment at scale.
Time Consumption Proportion. To better understand UniSage’s internal overhead, we divide its runtime into three stages: log sampling, trace sampling, and RCA. As shown in Figure 7(b), RCA accounts for 49% of the total time, trace sampling for 47%, and log sampling for only 4%. This breakdown demonstrates that our RCA module is relatively lightweight, despite being executed before sampling. It effectively addresses a key concern of the analysis-before-sample paradigm, the risk of high analysis latency hindering practical deployment. In contrast, UniSage maintains a balanced runtime profile suitable for real-world usage.
6. Discussion
6.1. Limitations
Most recent research on trace sampling, notably HindSight (hindsight, ), introduces the concept of retroactive sampling, where trace data is lazily retrieved to perform biased sampling earlier in the trace lifecycle. This approach improves the efficiency of biased sampling by deferring decisions until more information is available. While HindSight shares a similar motivation with UniSage (i.e. performing analysis before sampling), it focuses solely on edge-case traces and does not support downstream diagnosis tasks. In contrast, UniSage not only considers edge-cases but also emphasizes anomaly relevance and diagnosis support, enabling the retention of more informative and actionable data. Another limitation lies in the reliance on lightweight anomaly detection and RCA algorithms in the initial analysis stage. While these components are efficient and scalable, their accuracy may impact the quality of sampling decisions. Future work could investigate how more advanced or adaptive analysis techniques affect the sampling outcome.
6.2. Generalizability
While our current evaluation focuses on AD and RCA, we believe UniSage can generalize well to other downstream tasks such as performance profiling and workload characterization. For performance profiling, UniSage’s sampling strategy incorporates performance indicators (e.g., trace durations, metrics), which are commonly used in profiling tasks. This enables the retained data to still reflect key performance behaviors, even under storage constraints. For workload characterization, which aims to comprehensively identify workload patterns, UniSage is explicitly designed to capture diverse and rare behaviors (e.g., anomalies, edge cases). This makes it particularly suitable for characterizing workloads when observability data is limited. While our current evaluation focuses on AD and RCA, we agree that extending empirical validation to other tasks is valuable and plan to explore this in future work.
6.3. Threats to Validity
The main threat to internal validity lies in the implementation of baselines. For Sieve and TraStrainer, we use their publicly available open-source implementations. However, Sifter lacks open-source code, so we carefully re-implement it based on descriptions in the original papers, ensuring the use of identical libraries and settings whenever possible.
The primary threat to external validity concerns the generalizability of our results. Although we evaluate UniSage on two widely-used microservice benchmarks, the number of injected fault types remains limited (five in total), which may not capture the full diversity of real-world failure modes. Additionally, while some production datasets exist, they often include only traces without corresponding logs or metrics, preventing a comprehensive evaluation of our unified sampling strategy. To further validate the robustness of UniSage, future work could explore its application in more diverse system architectures, real production environments, and under more complex fault scenarios.
7. Related Works
Distributed tracing is a fundamental technique for monitoring and diagnosing the behavior of distributed systems. It captures the end-to-end execution of requests across services, providing visibility into performance bottlenecks and failure points. This is essential for identifying issues that are nondeterministic or difficult to reproduce in isolation. By tracing the flow of a request through different components, developers can better understand and debug complex distributed systems. Over the years, significant progress has been made in tracing frameworks. X-Trace (X-Trace, ) focuses on tracing across multiple system layers, enabling detailed visibility into application behavior. Dapper (dapper, ), Google’s tracing system, demonstrates the viability of large-scale distributed tracing in production environments. Canopy (canopy, ), developed at Facebook, captures performance data and causal relationships across entire request paths to support end-to-end analysis. In recent years, many open-source tracing frameworks such as Jaeger (Jaeger, ) and Zipkin (Zipkin, ) have been widely adopted in practice. Distributed tracing has enabled a wide range of system observability and analysis tasks, including performance profiling (mi2013toward, ; sambasivan2011diagnosing, ), anomaly detection (microhecl, ; tracecrl, ; nedelkoski2019anomaly, ), and failure diagnosis (gan2021sage, ; Eadro, ; Nezha, ; MRCA, ; li2025coca, ; liu2020unsupervised, ; hemirca, ; wu2025trace, ). As trace volume grows rapidly in production environments, several trace reduction techniques have been proposed (tracemesh, ; steam, ; huang2021sieve, ; sifter, ), with head sampling (canopy, ; dapper, ) and tail sampling (huang2021sieve, ; sifter, ; las2018weighted, ) being the most commonly adopted strategies.
Recent studies have started to incorporate troubleshooting and anomaly-awareness into sampling decisions. For example, TraStrainer (TraStrainer, ) introduces a trace sampling approach that adapts dynamically to system runtime state by leveraging metrics. By jointly considering trace diversity and metric anomalies, it prioritizes performance-relevant traces. However, TraStrainer’s detection scope is limited to anomalies that manifest in metric signals, leaving out other types of failures, and it only reduces traces. Hindsight (hindsight, ) presents a retroactive sampling framework targeting edge-case request diagnostics. Although it enhances sampling precision by using full-trace hindsight, it remains focused solely on traces and request-level analysis. In contrast, our work proposes a unified post-detection sampling framework that supports both traces and logs. By incorporating signals from metrics, logs, and traces, our system can detect and retain a broader range of anomaly types. Furthermore, our method incorporate interpretable signals that explain the rationale behind each sampling decision, revealing why the retained data is anomaly-related or represents uncommon behavior. This design effectively bridges the gap between fault diagnosis and data reduction.
8. Conclusion
In this paper, we presented UniSage, a novel post-analysis-aware sampling framework for observability data in distributed systems. Unlike traditional approaches that sample data prior to analysis, UniSage reverses this process by first performing lightweight anomaly detection and root cause analysis, and then retaining a concise yet informative subset of data for future use. Our method jointly considers both traces and logs, and emphasizes anomaly relevance and data diversity, ensuring that the preserved samples are not only valuable for fault diagnosis but also interpretable. Moreover, by addressing the previously overlooked problem of log sampling and offering a unified solution, UniSage fills an important gap in the observability pipeline. Extensive experiments on two widely-used microservice benchmarks demonstrate that UniSage outperforms existing sampling strategies in terms of effectiveness, efficiency, and support for root cause analysis. In future work, we plan to explore adaptive sampling strategies that dynamically adjust retention policies based on evolving system behavior, as well as extend our framework to support other modalities such as events and alerts.